# Neuromorphic Intermediate Representation: A Unified Instruction Set for Interoperable Brain-Inspired Computing

### v1

@@ Abstract
Spiking neural networks and neuromorphic hardware platforms that emulate neural dynamics are slowly gaining momentum and entering main-stream usage. Despite a well-established mathematical foundation for neural dynamics, the implementation details vary greatly across different platforms. Correspondingly, there are a plethora of software and hardware implementations with their own unique technology stacks. Consequently, neuromorphic systems typically diverge from the expected computational model, which challenges the reproducibility and reliability across platforms. Additionally, most neuromorphic hardware is limited by its access via a single software frameworks with a limited set of training procedures. Here, we establish a common reference-frame for computations in neuromorphic systems, dubbed the Neuromorphic Intermediate Representation (NIR). NIR defines a set of computational primitives as idealized continuous-time hybrid systems that can be composed into graphs and mapped to and from various neuromorphic technology stacks. By abstracting away assumptions around discretization and hardware constraints, NIR faithfully captures the fundamental computation, while simultaneously exposing the exact differences between the evaluated implementation and the idealized mathematical formalism. We reproduce three NIR graphs across 7 neuromorphic simulators and 4 hardware platforms, demonstrating support for an unprecedented number of neuromorphic systems. With NIR, we decouple the evolution of neuromorphic hardware and software, ultimately increasing the interoperability between platforms and improving accessibility to neuromorphic technologies. We believe that NIR is an important step towards the continued study of brain-inspired hardware and bottom-up approaches aimed at an improved understanding of the computational underpinnings of nervous systems. NIR is available at https://github.com/neuromorphs/NIR
@@ 1 Introduction
The human brain is an exemplar of computational efficiency, outperforming today’s advanced machine learning models and hardware in numerous measures, particularly regarding energy consumption. State-of-the-art machine learning algorithms require vast amounts of energy and computational resources, especially when trained on large datasets, whereas the brain operates on a tiny power budget, performing a myriad of complex tasks seamlessly. Unlike digital computations that rely on precise binary logic, the brain’s computations are predominantly analog in nature, utilizing gradients of ion concentrations and single-bit activations, commonly referred to as spikes. The biological circuits in the brain can be described as dynamical systems that evolve over time and are inherently noisy, yet they manage to process information with remarkable accuracy and adaptability [1, 2, 3]. Neuromorphic computing, defined as “circuits that emulate the temporal processing of signals in the brain” [2, p. 361], draws inspiration from the efficiency of the nervous systems to rethink the principles of information processing. As such, neuromorphic computing stands in contrast to the conventional and widespread von Neumann architecture, relying on entirely new algorithms, software tools and hardware [4].
Figure 1: High-level overview of what NIR is. (a) NIR allows for continuous-time representation of nodes that can then be executed on continuous-time hardware or simulators, or discretized for use on discrete-time hardware or simulators. (b) A taxonomy of discrete and continuous time hardware and simulators.
@@
Neuromorphic systems are typically approached bottom-up or top-down [5]. Rapid scientific progress in both directions means that neuromorphic platforms need to compete on both efficient and flexibility. Secondly, contemporary neuromorphic hardware and software are often developed together, similar to the biological system from which we draw inspiration—evidently a product of conjoined development between physical form and behavioral function [2]. blurring the lines between top-down function and bottom-up computation. The blurring of top-down and bottom-up approaches makes it difficult to extrapolate clear abstractions that generalize to other technology stacks, which impedes incremental technological progress. As a result, we face a landscape of heterogeneous neuromorphic simulators and design tools. By identifying domain-specific computational primitives, we argue that a layer of abstraction can be established to achieve greater interoperability and, in turn: (1) allow hardware and software efforts to develop independently and more rapidly, (2) lower the barrier-of-entry for newcomers, and (3) provide theoreticians with a more stable representation to reason about neuromorphic circuits.
@@
The field of deep learning faced similar incompatibility challenges a decade ago, but since then, intermediate representations and compiler frameworks such as ONNX [6], MLIR [7], XLA [8], and TVM [9] have bridged the gap between model description and different hardware accelerators. Specifically, these representations map models written in higher-level languages to different classes of hardware backends, such as CPUs, GPUs, TPUs, and FPGAs. Given the nature of clocked computation in the traditional computing stack, deep learning representations are based on digital instructions in discrete time. This contrasts our definition of neuromorphic models that are best captured as dynamical systems in continuous time. Any representation of a continuous system as a sequence of digital instructions implies fusing it with numerical integration techniques, thus intertwining the integration with the desired dynamics. Additionally, neuromorphic simulators and hardware platforms operate on the level of neuronal dynamics, a fundamentally different abstraction than digital intermediate representations for conventional machine instruction set architectures. This renders existing intermediate representations such as ONNX and MLIR poorly suited to describe neuromorphic computations faithfully.
In a joint effort between academia and industry, we present the Neuromorphic Intermediate Representation (NIR): an abstraction layer that bridges neuromorphic software and hardware platforms. NIR currently links 7 neuromorphic simulators (Lava, [10], Nengo [11], Norse [12], Rockpool [13], Sinabs [14], snnTorch [15], and Spyx [16]) and 4 neuromorphic hardware platforms (Loihi 2 [17] via Lava, Speck via Sinabs, SpiNNaker2 [18], and Xylo [19] via Rockpool). NIR defines graphs of computational nodes, where each node represents a computational primitive defined as a hybrid continuous-time dynamical system. This idealized description provides three distinct advantages compared to digital representations: (1) it avoids any assumptions around discretization or hardware constraints, (2) it provides a normative description against which implementations can be compared, and (3) it paves the way for seamless integration with mixed-signal chips, as well as analog accelerators and substrates. The computational graph representation is shown in Figure 1 along with its relation to digital and analog hardware as well as software, independently of the underlying continuous-time or discrete-time execution model.
In this paper, we (1) derive and implement a set of computational primitives that are common to the varied neuromorphic software and hardware landscape, (2) provide support for NIR across 7 different simulators and 4 different hardware platforms, (3) demonstrate flexible cross-platform deployment of NIR models, and (4) evaluate exemplary neuromorphic models defined in NIR across all 11 supported platforms—hardware and software. We develop our models on software simulators that are external and unfamiliar to the various neuromorphic hardware platforms we later execute the models on. The deployment of software-generated models on neuromorphic hardware was previously only possible through hardware-specific software libraries, and NIR effectively decouples the development of neural models from platform-specific tools. NIR now allows users to leverage tools and resources from the active neuromorphic open-source community and allows model development and platform-specific tools to develop efficiently in parallel, as illustrated in Figure 2). In this way, NIR increases accessibility to—and flexibility of—neuromorphic hardware platforms. This work comes at a decisive moment for novel compute technologies, where the power consumption of neural networks grows at an unprecedented rate, but also at a time when the platforms become available on a consumer scale for the first time.
1.1 Related work
Several works have called for widening usability and access to neuromorphic hardware and simulators [4, 20, 21] through co-design [5, 22] and shared representations [23]. PyNN [24] allows users to define neuroscientific spiking neural network (SNN) models in a simulator-independent language. PyNN models can be run without modifications on different neural simulators (NEST, Neuron, Arbor, Brian2) or neuromorphic hardware (e.g., BrainScaleS or SpiNNaker). PyNN is a programmatic frontend and does not provide a serializable data format. This is addressed by NeuroML [25] which supplies standards for defining and exchanging biological cell and network models. Both PyNN and NeuroML focus on computational neuroscience and not on the scalable compute capabilities of SNNs.
@@ 1.1 Related work
Several frameworks have been developed to design neuromorphic algorithms. Fugu [26] provides a high-level API to define and combine spiking neural algorithms. It generates an SNN graph as an intermediate representation which can serve as input to neuromorphic simulators or hardware compilers. Similarly, Lava [10] is a framework for developing brain-inspired neural network models and mapping them to Intel’s digital neuromorphic hardware. Its intermediate representation implements the Communicating Sequential Processes (CSP) principle [27], where neurons are represented by processes which send spikes via channels connected to other processes. Lava significantly simplifies the programming of the neuromorphic chip Loihi 2 [28]. However, at its core, it requires synchronization messages for triggering neuron updates, which is incompatible with analog neuromorphic systems. In addition to Fugu and Lava, there are frameworks such as the SNN-Toolbox [29] or NxTF [30] for converting neural network models from ML frameworks to specific SNN simulators and neuromorphic hardware. Zhang et al. [31] have developed an abstraction hierachy for brain-inspired computing with loose guarantees, portable across both von Neumann and neuromorphic architectures. Other interfaces focus on particular computing elements, such as crossbar arrays using beyond-CMOS technologies [32, 23, 33], or ReLU and leaky integrate-and-fire units [34]. Unfortunately, none of these standards have spread beyond their intended subdomain within neuromorphic computing.
@@
Figure 2: a-d) shows four NIR primitives, where the name of the primitive is highlighted on top, the implemented computation is illustrated in the white box, and the parameters that are stored with the NIR graph are highlighted on the bottom. a) shows a stateful primitive, while b) and c) show stateless primitives, and d) shows a higher-order primitive. e-f) illustrate the concept of an intermediate representation (IR) with 7 software (SW) and 4 hardware (HW) backends. Instead 30 different compilers covering all $m\times n$ cases (e)), only $m+n$ interfaces between each hardware and software platform to NIR are necessary (f)).
@@ 2 Results
The Neuromorphic Intermediate Representation (NIR) uses signal-flow graphs to describe the intricate dynamics of neuromorphic neural networks. NIR captures coupled hybrid systems in a structured graph representation, shown in Figure 1 and 2, which are formally introduced in Section 2.2. In contrast to existing representations of neural dynamics, NIR operates on a predefined set of computational primitives rather than custom neural equations, which cannot easily be supported on efficient and specialized hardware. This approach is motivated by the nearly universal agreement on a few neuron models, such as the leaky integrate-and-fire model, in prevalent neuromorphic hardware platforms and simulators. The reductionist approach allows us to achieve strong generality early on; NIR is presently supported by four digital neuromorphic hardware platforms: Intel Loihi 2 [10], SynSense Speck, SpiNNaker2 [18], and SynSense Xylo [19]. Additionally, 7 software simulators now support NIR: Lava and Lava-DL, [10], Nengo [11], Norse [12], Rockpool [13], Sinabs [14], snnTorch [15], and Spyx [16].
@@
The primary strength of NIR resides in its ability to reproduce the same computation on heterogeneous platforms, abstracting away platform-specific intricacies. Below, we demonstrate how a NIR graph can be executed and supported by the above-mentioned platforms across three heterogeneous tasks: a leaky integrate-and-fire model, a spiking convolutional network, and a recurrent reservoir network. Later, in Sections 2.2 and 2.3, we establish the axioms underpinning NIR graphs and list the computational primitives we presently support.
It is important to note that the hardware constraints and implementation choices of each platform necessarily lead to divergent behaviors. NIR does not attempt to match the computation perfectly on every platform. Rather, NIR aims to establish an idealized model that serves as a common ground for comparative analysis while highlighting the inter-platform discrepancies. By disentangling model dynamics from hardware idiosyncrasies, NIR encourages the development of platform-independent models that can subsequently be tuned to a specific platform, using, for instance, quantization-aware training. Our contribution serves a broader objective of neuromorphic computing: to simplify the creation of biologically inspired efficient neural networks, facilitating their consistent functionality across a wide range of computational environments.
2.1 Experiments
@@
We conduct a series of experiments spanning diverse neuromorphic platforms to gauge the robustness and versatility of NIR. We chose three distinct tasks that are common within neuromorphic computing, through which we assess and compare performance across all compatible platforms: a leaky integrate-and fire neuron, a spiking convolutional neural network, and a recurrent reservoir model (see Figure 3).
@@
Figure 3: The experiments we evaluate across 11 neuromorphic platforms and example data from the NMNIST and Braille dataset. (a) A single leaky integrate-and-fire neuron model, (b) a recurrent Braille classification model using current-based leaky integrate-and-fire in a spiking recurrent neural network (SRNN), and (c) a spiking convolutional neural network (SCNN). (d) and (e) shows sample data for the NMNIST and Braille datasets, respectively. The NMNIST activity is averaged across 300 timesteps.
@@
A central difficulty in transferring a computation from one neuromorphic platform to another is that the neuron models and discretization schemes do not always match up exactly. There is a large amount of variation between differently implemented neuron models, even for the seemingly simple leaky integrate-and-fire (LIF) neuron. A more detailed discussion of the subtle differences in neuron models supported by NIR is presented in Section 4.4.
2.1.1 Leaky integrate-and-fire dynamics
@@
As a first experiment, we demonstrate the behavior of a single leaky integrate-and-fire neuron under an identical input spike train (Figure 4 (a)) for every available platform. The LIF neuron was first instantiated on the software platform Norse, then exported into a NIR graph, and subsequently imported and executed by all different platforms. Figure 4 (a) visualizes both the voltage traces and output spikes from the platforms.
@@
Figure 4: Experimental results. (a) Single leaky integrate-and fire neuron ordered from top to bottom: input spikes, voltage traces, and output activations. The timestep of spikes are well-aligned and only differ systematically for Rockpool and Sinabs due to discretization differences. Membrane potentials are normalized to lie between the resting potential $v_{rest}=0$ and the firing threshold $\vartheta$ to ignore platform-specific details regarding the numerical representations. (b) Spiking convolutional neural network: a platform-by-platform comparison of the spiking activity from the first spiking layer using cosine similarity (1 equals perfect overlap). Sinabs, snnTorch and Speck deviate from most other implementations due to their discretization choices. (c) and (d) Spiking recurrent neural network: similarity measure between the spiking activity of the first CuBa-LIF layer for an SRNN with biases and reset to zero and an SRNN without biases and subtractive reset.
@@
${}^{\dagger}$ Nengo Norse Rockpool Sinabs snnTorch Speck SpiNNaker2 Spyx Xylo
SCNN 98.2% 98.1% 98.1% N/A 98.5% 97.9% 95.4% 98.2% 97.1% N/A
SRNN (Subtr.) N/A 78.6% 93.57% 71.4% N/A 92.1% N/A 93.57% 92.12% 85.71%
SRNN (Zero) 48.6% 55.7% 94.29% N/R N/A 95.0% N/A 85.00% 84.29% N/A
@@
Table 1: Experimental accuracies for the spiking convolutional neural network (SCNN) and the spiking recurrent neural network (SRNN) measured against unseen test dataset. ${}^{\dagger}$ Lava offers a bitwise accurate simulator for Loihi 2, which is why the Loihi 2 column is merged with Lava.
@@
To assess the viability of NIR as an intermediate representation for neuromorphic computing, we ensure that the dynamics are consistent across all platforms regarding the most common ways of encoding information in spiking neural networks—namely, rate encoding and spike time encoding [15]. The equivalence of output firing rates across different platforms can easily be seen, as the number of output spikes is identical across all platforms and since the spiking timings mostly coincide. Rockpool and Sinabs notably differ because of their integration scheme. They integrate the incoming signal ($v^{-}$, equation 1a) when checking for a threshold crossing before they update the leak term ($v^{+}$, equation 1b), whereas most other simulators immediately integrate the leak (equation 2). This effectively brings the neuron closer to the firing threshold earlier on, as observed in Figure 4 (a).
$v^{-}(t+1)$ $=\frac{dt}{\tau}\left(v(t)+v_{\text{leak}}+Ri(t+1)\right)$ (1a)
$v^{+}(t+1)$ $=v(t)-v^{-}(t+1)$ (1b)
@@
We further observe that the spike times are systematically delayed by one timestep for Lava and Loihi 2. This is because Lava calculates whether a neuron spikes at timestep $t$ based on its membrane potential at timestep $t-1$, i.e., before the addition of the input to the membrane, rather than after the input was added to the membrane potential. This systematic difference will only have an effect if absolute spike timing is used, but will have no impact if relative spike timing is used, that is if information is encoded in the timing between spikes.

### v2

@@ Abstract
Spiking neural networks and neuromorphic hardware platforms that simulate neuronal dynamics are getting wide attention and are being applied to many relevant problems using Machine Learning. Despite a well-established mathematical foundation for neural dynamics, there exists numerous software and hardware solutions and stacks whose variability makes it difficult to reproduce findings. Here, we establish a common reference frame for computations in digital neuromorphic systems, titled Neuromorphic Intermediate Representation (NIR). NIR defines a set of computational and composable model primitives as hybrid systems combining continuous-time dynamics and discrete events. By abstracting away assumptions around discretization and hardware constraints, NIR faithfully captures the computational model, while bridging differences between the evaluated implementation and the underlying mathematical formalism. NIR supports an unprecedented number of neuromorphic systems, which we demonstrate by reproducing three spiking neural network models of different complexity across 7 neuromorphic simulators and 4 digital hardware platforms. NIR decouples the development of neuromorphic hardware and software, enabling interoperability between platforms and improving accessibility to multiple neuromorphic technologies. We believe that NIR is a key next step in brain-inspired hardware-software co-evolution, enabling research towards the implementation of energy efficient computational principles of nervous systems. NIR is available at neuroir.org
@@ 1 Introduction
The human brain is an exemplar of computational efficiency, outperforming today’s advanced machine learning models and hardware in multiple measures, particularly regarding energy consumption. State-of-the-art machine learning algorithms require vast amounts of energy and computational resources, especially when trained on large datasets [1, 2] whereas the brain operates on a tiny power budget, learning and performing a myriad of complex tasks seamlessly. Unlike digital computations that rely on precise binary logic, the brain’s computations have a significant analog component in nature, utilizing gradients of ion concentrations and action potentials, which can be represented as unary activations, commonly referred to as spikes. Neuromorphic computing, defined as “circuits that emulate the temporal processing of signals in the brain” [3, p. 361], draws inspiration from the efficiency of the nervous systems to rethink the principles of information processing. As such, neuromorphic computing stands in contrast to the conventional and widespread von Neumann architecture, and relies on entirely new algorithms, software tools and hardware [4].
Figure 1: High-level overview of NIR. (a) NIR allows for continuous-time representation of specific models that can then be executed on continuous-time hardware or simulators, or discretized for use on discrete-time hardware or simulators. (b) A taxonomy of discrete and continuous time hardware and simulators. Some representative hardware systems are shown. Analog Continuous Time: BrainDrop [5], DYNAP-SE [6], and emerging technologies. These systems are greyed out because they are not covered in the present paper. Analog Discrete Time: NeuRRAM [7], UMich RRAM-IMC [8]. Digital Continuous Time: Akida, SpiNNaker 1/2 [9, 10], Speck. Note that we include chips with asynchronous routing in this category. Digital Discrete Time: Generic FPGA-based SNN Accelerators, ReckOn [11], Xylo [12].
@@
Much like how the brain can be thought of as a physical embodiment of the neural code consisting of architecture, hardware, and software, contemporary neuromorphic hardware and software are often developed together [3]. The blurring of top-down and bottom-up approaches makes it difficult to extrapolate clear abstractions that generalize to other technology stacks, which impedes incremental technological progress [13]. The blurring of top-down and bottom-up approaches makes it difficult to extrapolate clear abstractions that generalize to other technology stacks, which impedes incremental technological progress. As a result, we face a landscape of heterogeneous neuromorphic simulators and design tools. By identifying domain-specific computational primitives, we argue that a layer of abstraction can be established to achieve greater interoperability and, in turn: (1) allow hardware and software efforts to develop independently and more rapidly, (2) lower the barrier-of-entry for newcomers, (3) provide theoreticians with a stable representation to study fundamentals of computing across different neuromorphic circuits, and (4) possibly enable more energy-efficient computing.
@@
The field of deep learning faced similar heterogeneity and incompatibility challenges a decade ago, which have been addressed by intermediate representations (IR) and compiler frameworks such as ONNX [14], MLIR [15], XLA [16], and TVM [17]. Note that, by definition, IRs serve exclusively as a model description, while the compilers carry out the translation between platforms, although they are often mixed in practice. These tools have reduced the gap between the simulation of an application and different hardware accelerators, enabling models written in higher-level languages to be mapped to different classes of hardware backends, such as CPUs, GPUs, TPUs, and FPGAs. Notably, the development of sophisticated compiler frameworks was only possible once a stable representation and exchange format were established. Following this precedent, NIR aims to establish a stable representation and exchange format for neuromorphic computing, as a foundation on which more specialized compiler frameworks can be built. Similarly, cross-framework compatibility for deep learning has been addressed by approaches such as Ivy [18] or MMdnn [19]. Given the nature of clocked computation in the traditional computing stack, deep learning representations are based on digital instructions in discrete time. This contrasts our definition of neuromorphic models that are best captured as continuous-time dynamical systems. Any representation of a continuous system as a sequence of digital instructions implies using numerical integration techniques, possibly introducing numerical errors during computation of the desired dynamics. Additionally, neuromorphic simulators and hardware platforms operate on the level of neuronal dynamics, a fundamentally different abstraction than digital intermediate representations for conventional machine instruction set-based architectures. This renders existing intermediate representations such as ONNX and MLIR suboptimal for describing neuromorphic computations.
Numerous configurable neuromorphic systems have been developed over the last decades [20]: the implementation approaches range from analog / mixed-signal [21, 22, 5, 23], over digital [24, 25, 26], hybrid analog/digital [6] to processor-based [9, 10] solutions. We refer to Furber 2016 [27] and Thakur et al. 2018 [28] for a detailed overview and comparison of the most prominent large-scale systems. Most of the systems offer custom interfaces for programming the chips, such as the Corelet approach [29] for TrueNorth [24] or PyNCS [30] for the UZH—ETHZ chips [6].
Several works emphasize the need to make neuromorphic hardware and simulators more user-friendly and accessible [4, 31, 32] through co-design [13] and shared representations [33]. PyNN [34] is a Python library that allows users to define spiking neural network (SNN) models in a simulator-independent language. PyNN models can be run without modifications on different neural simulators (NEST [35], Neuron [36], Arbor [37], Brian2 [38], CARLSim [39]). PyNN has developed into the most wide-spread common interface to neuromorphic hardware, supporting the Heidelberg Spikey chip [40], the BrainScaleS systems [41] and SpiNNaker1 [42]. NeuroML [43] provides a serializable set of biological cell and network models which emphasizes computational correctness for neuroscientific studies. The popular Brian2 simulator [38] has been extended to interface other tools such as the SNN-GPU simulator GeNN [44] or to emulate the Intel Loihi chip [45].
@@
Several frameworks have been developed to design neuromorphic algorithms. Fugu [46] provides a high-level API to define and combine spiking neural algorithms. It generates an SNN graph as an intermediate representation, which can serve as input to neuromorphic simulators or hardware compilers. Similarly, Lava [47] is a framework for developing brain-inspired neural network models and mapping them to Intel’s digital neuromorphic hardware. Its representation is based on the Communicating Sequential Processes (CSP) principle [48], where neurons are represented by processes which send spikes via channels connected to other processes. Lava significantly simplifies the programming of the neuromorphic chip Loihi 2 [49]. However, at its core, it requires synchronization messages for triggering neuron updates, which is incompatible with applications exhibiting continuous-time behaviors. Nengo [50] is a mature library for brain simulation and deep learning inspired spiking networks. Similar to PyNN, it has been connected to several neuromorphic systems such as BrainDrop [5], Loihi, SpiNNaker1 and FPGAs. In addition to Fugu, Lava and Nengo, there are frameworks such as the SNN-Toolbox [51], NxTF [52], or hxtorch.snn [53] for converting neural network models from ML frameworks to specific SNN simulators and neuromorphic hardware. Zhang et al. [54] have developed an abstraction hierarchy for brain-inspired computing with loose guarantees, portable across both von Neumann and neuromorphic architectures. Other interfaces focus on particular computing elements, such as crossbar arrays using beyond-CMOS technologies [55, 33], or ReLU and leaky integrate-and-fire units [56]. Unfortunately, none of these standards have spread beyond their intended subdomain within neuromorphic computing.
Theoretically, our work is inspired by the work on signal flow graphs [57], linear time-invariant systems [58], and mealy machines [59] for the description of composable circuit components with applications to digital, or analogue, or hybrid systems. Compiler infrastructures projects like LLVM [60], and later MLIR [15], has been leading the field in optimizing and transforming code for a wide range of programs and architectures.
In this work, we (1) derive and implement a set of model-centric computational primitives that are common to multiple neuromorphic software and hardware systems, (2) illustrate the application of NIR across 7 different simulators and 4 different hardware platforms, (3) demonstrate flexible cross-platform deployment of NIR models, and (4) evaluate exemplary neuromorphic models defined in NIR across all 11 supported platforms—in hardware and software. This work comes at a decisive moment for novel compute technologies, where the power consumption of neural networks is growing at unprecedented rates, and when neuromorphic platforms are available at scale to the consumer for the first time.
@@
Figure 2: Composition of primitives and their mapping to and from software and hardware systems. (a-d) shows four NIR primitives, where the name of the primitive is highlighted on top, the implemented computation is illustrated in the white box, and the parameters that are stored with the NIR graph are highlighted on the bottom. (a) shows a stateful primitive, while (b) and (c) show stateless primitives, and (d) shows a higher-order primitive. (e-f) illustrate the concept of an intermediate representation (IR) with 7 software (SW) and 4 hardware (HW) backends. Instead of 30 different compilers covering all $m\times n$ cases (e), only $m+n$ interfaces between each hardware and software platform to NIR are necessary (f).
@@ 2 Results
In a collaboration between academia and industry, we propose and illustrate the Neuromorphic Intermediate Representation: a model-centric abstraction layer that simplifies the translation between simulated applications, neuromorphic software, and digital hardware platforms. NIR currently links 7 neuromorphic simulators (Lava [47], Nengo [50], Norse [61], Rockpool [62], Sinabs [63], snnTorch [64], and Spyx [65]) and 4 digital neuromorphic hardware platforms (Loihi 2 [66] via Lava, Speck via Sinabs, SpiNNaker2 [67], and Xylo [68] via Rockpool). NIR represents computations as graphs, where each node represents a computational primitive defined by a hybrid continuous-time dynamical system. This idealized description provides three distinct advantages: (1) it avoids any assumptions around discretization or hardware constraints, (2) it provides a reference model against which implementations can be compared, and (3) it decouples the software description from the hardware layer, allowing seamless integration with mixed-signal chips, as well as analog hardware, and hybrid digital-analog systems. The computational graph representation is shown in Figure 1 along with its relation to digital and analog hardware as well as software, independently of the underlying execution model being continuous-time or discrete-time. While the continuous nature of NIR can be applied to mixed-signal hardware platforms, this work focuses on digital simulators and hardware platforms. It is also important to point out that the conversion between continuous-time dynamical systems and digital platforms will give divergent results. We will further define and measure these discrepancies in later sections.
@@
NIR defines a set of primitives as coupled hybrid systems, formally introduced in the Methods Section, that can be arbitrarily composed as shown in Figures 1 and 2. NIR operates on a predefined set of computational primitives rather than custom neural equations, which cannot easily be supported on efficient and specialized hardware. This approach is motivated by the nearly universal agreement on a few computational models, such as the leaky integrator and the integrate-and-fire models, in prevalent neuromorphic hardware platforms and simulators. It also allows us to achieve strong generality early on; NIR is presently supported by four digital neuromorphic hardware platforms: Intel Loihi 2 [47], SynSense Speck, SpiNNaker2 [67], and SynSense Xylo [68]. Additionally, 7 software simulators now support NIR: Lava (and Lava-DL) [47], Nengo [50], Norse [61], Rockpool [62], Sinabs [63], snnTorch [64], and Spyx [65].
@@
The primary strength of NIR resides in its ability to reproduce the same computation on heterogeneous platforms, abstracting away platform-specific differences. Thus, NIR is able to translate a computational model into multiple, and possibly different, hardware platforms; training or defining models can be done independent of the substrate carrying out the computation. Below, in Section Experiments, we demonstrate how a NIR graph can be executed and supported by the above-mentioned platforms for three different tasks: a leaky integrate-and-fire model, a spiking convolutional network, and a spiking recurrent neural network. Later, in Sections NIR axioms and Computational primitives, we establish the axioms underpinning NIR graphs and list the computational primitives we presently support.
Experiments
@@
We conduct a series of experiments spanning diverse neuromorphic platforms to gauge the robustness and versatility of NIR. We chose three distinct tasks that are common within neuromorphic computing, through which we assess and compare performance across all compatible platforms: a single leaky integrate-and fire (LIF) neuron, a spiking convolutional neural network (SCNN), and a spiking recurrent neural network (SRNN) (see Figure 3).
The single LIF neuron task is a simple task which allows us to assess the visual similarity of the computations on different simulator and hardware platforms. The SCNN and SRNN represent widespread types of neural networks. Whereas CNN-based networks are very common for visual data, RNN show their strengths in tasks involving temporal relations with a limited number of input dimensions.
@@
Figure 3: Computational graphs and sample data used in the experiments. (a) A single leaky integrate-and-fire neuron model (LIF), (b) a recurrent Braille classification model using current-based leaky integrate-and-fire (CuBa-LIF) in a spiking recurrent neural network (SRNN), and (c) a spiking convolutional neural network (SCNN). (d) and (e) shows sample data for the N-MNIST and Braille datasets, respectively. The N-MNIST activity is averaged across 300 timesteps.
@@
A central difficulty in transferring a computation from one neuromorphic platform to another is that the neuron models and discretization schemes do not always match up exactly. There will be variation between differently implemented neuron models, even for the seemingly simple LIF neuron. We refer to Section Discussion of mismatches for a more detailed discussion of the experimental mismatches, and the Supplementary Material, Section A, for the subtle differences in neuron models supported by NIR.
Leaky integrate-and-fire dynamics
@@
As a first experiment, we demonstrate the behavior of a single leaky integrate-and-fire neuron under an identical input spike train (Figure 4 (a)) for every available platform. The LIF neuron was first instantiated in Norse, then exported into a NIR graph, and subsequently imported and executed by all different platforms. Figure 4 (a) visualizes both the recorded voltage traces and output spikes from the platforms.
@@
Figure 4: Experimental results. (a) Single leaky integrate-and fire neuron ordered from top to bottom on different platforms: input spikes, voltage traces, and output activations. The timestep of spikes are well-aligned and only differ systematically for Rockpool and Sinabs due to discretization differences. Membrane potentials are normalized to lie between the resting potential $v_{rest}=0$ and the firing threshold $\vartheta$ to ignore platform-specific details regarding the numerical representations. (b) Spiking convolutional neural network: a platform-by-platform comparison of the spiking activity from the first spiking layer using cosine similarity (1 equals perfect overlap). Sinabs, snnTorch and Speck deviate from most other implementations due to their discretization choices. (c) and (d) Spiking recurrent neural network: similarity measure between the spiking activity of the first CuBa-LIF layer for an SRNN with biases and reset to zero and an SRNN without biases and subtractive reset. See main text for details on the similarity metrics.
Simulator Both Hardware
@@
Nengo Norse Rockpool Sinabs snnTorch Spyx Lava Speck SpiNNaker2 Xylo
SCNN 98.1% 98.1% N/A 98.5% 97.9% 97.1% 98.2% 95.4% 98.2% N/A
SRNN (Subtr.) 78.6% 93.57% 71.4% N/A 92.1% 92.12% N/A N/A 93.57% 85.71%
SRNN (Zero) 55.7% 94.29% N/R N/A 95.0% 84.29% 48.6% N/A 85.00% N/A
@@
Table 1: Experimental accuracies. Performance for the spiking convolutional neural network (SCNN) and the spiking recurrent neural network (SRNN) measured against unseen test dataset. The platform on which the spiking neural network was originally trained on is underlined. The Lava column stands for both the Loihi 2 chip and the Lava simulator, since they are bitwise accurate.
@@
To assess the viability of NIR as an intermediate representation for neuromorphic computing, we ensure that the dynamics are consistent across all platforms regarding the most common ways of encoding information in spiking neural networks—namely, rate encoding and spike time encoding [64, 69]. The equivalence of output firing rates across different platforms can easily be seen, as the number of output spikes is identical across all platforms and since the spiking timings mostly coincide. Rockpool and Sinabs notably differ because of their integration scheme. They integrate the incoming signal ($v^{-}$, equation 1) when checking for a threshold crossing before they update the leak term ($v^{+}$, equation 2), whereas most other simulators immediately integrate the leak (equation 3). This effectively brings the neuron closer to the firing threshold earlier on, as observed in Figure 4 (a).
$v^{-}(t+1)$ $=\frac{dt}{\tau}\left(v(t)+v_{\text{leak}}+Ri(t+1)\right)$ (1)
$v^{+}(t+1)$ $=v(t)-v^{-}(t+1)$ (2)
@@
We further observe that the spike times are systematically delayed by one timestep for Lava and Loihi 2. This is because Lava calculates whether a neuron spikes at timestep $t$ based on its membrane potential at timestep $t-1$, i.e., before the addition of the input to the membrane, rather than after the input was added to the membrane potential. This systematic difference will only have an effect if absolute spike timing is used, instead of relative spike timing, that is, if information is encoded in the timing between spikes.